OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example 7(sqrt)a can be written as 7(sqrt)a.

OpenStudy (anonymous):

o_O What's the problem, exactly? This just looks like declarative information. To which I'm like,"Mmmmk..." Is there something you're asked to solve for or simplify?

OpenStudy (anonymous):

\[7\sqrt{a}\] and...?

OpenStudy (anonymous):

Noooo lol

OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely

OpenStudy (lalaly):

\[\sqrt{36a}=\sqrt{6}\sqrt{6}\sqrt{a}=6\sqrt{a}\]

OpenStudy (anonymous):

Thanks.